OpenStudy (anonymous):

Given:Triangle A B C is similar to triangle X Y Z~Triangle A B C is similar to triangle X Y Zwith altitudes CR and ZS. ZS = 12, XY = 15, YZ = 18 and AB = 20, find the measure of CR.

OpenStudy (anonymous):

use triangle similarity properties proportions are equal

OpenStudy (anonymous):

From the similarity we have AB/XY=BC/YZ=AC/XZ=CR/ZS=20/18=10/9

OpenStudy (anonymous):

we also know that ZS=12 so CR/12=10/9 then solve for CR

OpenStudy (anonymous):

oh sorry error AB/XY should be 20/15 not 20/18 20/15 is reduced to 4/3 so CR/12=4/3